market demand and supply equations for Plywood are given by
Qs = 20,000 + 30P
Qd = 40,000 – 20P
1.Determine the equilibrium price and quantity
The Plywood industry is perfectly competitive, and the marginal cost equation for one firm, Greenply,is given by MC = 200 + 4Q
2.What is the short-run output rate for Greenply?Average Cost is given by
AC= 1000/Q +200 + 2Q
4.In the short-run , how much economic profit will the firm earn?
1
Expert's answer
2019-05-08T09:50:10-0400

1. The equilibrium price and quantity are:

Qs = Qd,

20,000 + 30P = 40,000 – 20P,

50P = 20,000,

P = 400,

Q = 20,000 + 30P = 32,000 units.

2. If the Plywood industry is perfectly competitive, then it maximizes profits, when P = MC.

400 = 200 + 4Q,

Q = 50 units is the short-run output rate for Greenply.

4. If AC= 1000/Q +200 + 2Q, then economic profit of the firm is:

TP = (P - ATC)×Q = (400 - 1000/50 - 200 - 2×50)×50 = 4000.